OpenStudy (anonymous):

Jack usually mows his lawn in 6 hours. Marilyn can mow the same lawn in 4 hours. How much time would it take for them to mow the lawn together?

OpenStudy (amistre64):

how much do they mow in 1 hour? j = 1/6 ; m=1/4 in one hour they get 10/24 = 5/12 done 12/5 gets your total

OpenStudy (amistre64):

2.4 hours

OpenStudy (amistre64):

2 hr 24 mins ?perhaps?